A car with a mass 1078 kg and moving at a speed 5 m/s hits a motionless truck of mass 2998 kg whose brake is not engaged.

If the car comes to a complete stop during the collision, how fast does the truck start moving?
Answer #1

here,

the mass of car , m1 = 1078 kg

the initial velocity of car , u1 = 5 m/s

the mass of truck ,m2 = 2998 kg

the final velocity of car , v1 = 0 m/s

let the final velocity of truck be v2

using conservation of momentum

m1 * u1 + m2 * u2 = m1 * v1 + m2 * v2

1078 * 5 + 2998 * 0 = 1078 * 0 + 2998 * v2

solving for v2

v2 = 1.8 m/s

the final velocity of truck is 1.8 m/s